# Smoothing of Headland Path Edges and Headland-to-Mainfield Lane Transitions Based on a Spatial Domain Transformation and Linear Programming

### v3

@@ 3.3 Problem 1: Headland path smoothing
In a first step, the segment is scaled by heuristics and replaced by a 5-point piecewise-affine (PWA) path as shown in Fig. 4(a). Several comments are made. First, the key characteristic of this scaling is that the tip of the resulting 5-point PWA path shall ensure zero spraying gap. Thus, the distance between tip and its projection point along the field contour must be less or equal than half the machinery operating width $w>0$. Second, beyond this key characteristic the shaping of the reference path is a heuristic and free choice. Here, a 5-point PWA path segment is used because it rounds out the original edgy headland path segment (more than e.g. an alternative 3-point PWA path segment would) and thereby helps the downstream linear program formulation by providing an improved reference around which system dynamics are linerised and discretised. Third, the heuristic nature of this first reference path generation step is acknowledged and mentioned in the Conclusion as a potential avenue of future research. Fourth and as a detail, we do not use popular Chaikin’s algorithm (see Chaikin (1974)) for corner cutting and further smoothing of the reference path, since this would replace every edge point with two edge points at each iteration step. Furthermore, it would cut off a significant portion of the tip of the 5-point PWA line after the first iteration.
In a second step, using the 5-point PWA path from Step 1 as reference path the following basic LP is proposed:
@@
Steering actuation constraints are (6c) and (6d) as discussed in Sect. 3.1. In practice, constraints (6b) which enforce the resulting path to stay on one side of the reference path are relaxed to, $e_{y,j}-e_{y,j}^{\text{ref}}\leq\sigma$, by introducing a non-negative slack variable $\sigma\geq 0$ and adding it to the objective function 6a with a high weight such that $\sum\nolimits_{j=1}^{N}|e_{y,j}-e_{y,j}^{\text{ref}}|$ is replaced with $\sum\nolimits_{j=1}^{N}|e_{y,j}-e_{y,j}^{\text{ref}}|+10^{16}\sigma$. Furthermore, for the absolute value arguments $N$ non-negative surrogate variables are introduced and the set of inequality constraints is extended accordingly. In the end, there are $n_{u}=2N+1$ scalar real-valued optimisation variables. Typical sizes of the LP (i.e., the number of variables and inequality constraints) are stated in the numerical example Section 4.
For generality, reference $e_{y,j}^{\text{ref}}$ is included in (6). In practice, it is typically set to zero, $e_{y,j}^{\text{ref}}=0,\forall j=1,\dots,N$.
One benefit of proposed formulation (6) is that it is hyperparameter-free. There are no weighting parameters.
An even smaller alternative LP was considered, with a minmax-objective function, $\min\nolimits_{\{\delta_{j}\}_{j=0}^{N-1}}\ \max\nolimits_{j\in\{1,\dots,N\}}|e_{y,j}-e_{y,j}^{\text{ref}}|$, replacing (6a). However, this was found to be unsuited since only the maximum deviation from the reference path would be penalised, which resulted in jagged trajectories not accurately tracking the complete reference path.
After the solution of (6), a smoothed headland path segment as shown in Fig. 4(b) results.
Finally, LP-iterations or refinement steps can be conducted. Thus, for the first solution of (6) a 5-point-PWA path is used as reference. Then, the resulting path is used as reference for the second solution of (6). In theory, many of such LP-iterations can be carried out. In practice, one refinement step was found to be useful and sufficient.
@@ 3.4 Problem 2: Headland-to-mainfield lanes transitions
Before stating the LP for Problem 2, the term headland-to-mainfield lanes transitions is clarified. It is used comprehensively for brevity but shall always refer to both, (i) transitions from headland to a mainfield lane, but also (ii) transitions from a mainfield lane to the headland path. Both types of transitions naturally occur for a field coverage path. As will be shown below, a subtle difference in constraints formulation is applied to differentiate the two types.
@@
In practice, the objective function (7a) is slightly refined to better avoid undesired overshoots out of the mainfield area and into the headland area. Therefore, (7a) is replaced with $\sum\nolimits_{j=1}^{N}c_{j}|e_{y,j}-e_{y,j}^{\text{ref}}|$ and weights
@@
for transitions from mainfield lane to the headland path, where index $I$ is computed from the reference path as its last and first index within a 1m-radius to the headland path for the two transition types, respectively. The value 100 is selected exemplarily and can be any large constant, such that the objective function of the LP can still be considered hyperparameter-free. The benefit of this weighted approach is that no additional inequality constraints need to be introduced. As a result, LP (7) for Problem 2 is smaller than LP (6) for Problem 1.
As for above LP (6), for the absolute value arguments in (7a) $N$ non-negative surrogate variables are introduced and the set of inequality constraifnts is extended accordingly. In the end, there are $n_{u}=2N$ scalar real-valued optimisation variables. Typical sizes of the LP (i.e., the number of variables and inequality constraints) are stated in the numerical example Section 4. As in the previous subsection, reference $e_{y,j}^{\text{ref}}$ in (7) is typically set to zero, $e_{y,j}^{\text{ref}}=0,\forall j=1,\dots,N$.
Fig. 6 visualizes the result of a headland-to-mainfield lane transition. As part of Step 1 of the hierarchical algorithm for Problem 2 a Dubins path is fitted and replacing the edgy original transition. The Dubins path, plus an additional small path segment before and after selected based on heuristics to permit additional maneuvering space, serves as reference to Step 2 of the hierarchical algorithm based on which (7) is formulated and solved.
Fig. 6 is of special interest for two reasons. First, it illustrates how Dubins paths are not suited for the generation of precision paths to be followed by autonomous machines online, especially when calculating Dubins paths with minimal turning radius of the machine (see Fig. 6(a)). Due to actuation constraints deviations would be expected in any online tracking application. Second, it illustrates the effect on tracking performance when using different turning radii $R_{\text{Dubins}}$ in the Dubins reference path generation step. When using the minimal turning radius Fig. 6(a) results. By using an artificially larger $R_{\text{Dubins}}$ the LP-solution trajectory follows more closely the reference path, see Fig. 6(b) and 6(c). The drawback of a larger $R_{\text{Dubins}}$ is that the resulting path is longer and more compacted area occurs from corresponding tyre traces and the corresponding steering maneuvers, which reduces growable crop area in the field area. A trade-off with a preference towards small $R_{\text{Dubins}}$ for the reference path generation step is recommended.
Finally, in contrast to Problem 1 and its discussion in Section 3.3 for the case of Problem 2 no LP-iterations are conducted. It was empirically found that they did not improve results and are therefore omitted.
@@
Fig. 6: Problem 2: The effect on tracking performance when using three different $R_{\text{Dubins}}$ in the reference path generation step is visualised. The drawback of a larger $R_{\text{Dubins}}$ is that the resulting path is longer and more compacted field area occurs from corresponding tyre traces and steering maneuvers, which reduces growable crop area. A trade-off with tendency towards small $R_{\text{Dubins}}$ is recommended.
3.5 Discussion of Parameters and Practical Details
This section elaborates on important parameter choices and other practical details.
@@ 3.5 Discussion of Parameters and Practical Details
First, the selection of discretisation spacing, $D_{s,j}=s_{j+1}-s_{j},~{}\forall j=0,\dots,N-1$, is discussed, which typically is selected as uniformly spaced but can also be non-uniformly spaced. Its importance for spatial-based numerical optimisation is underlined. Its selection influences on one hand the size of the LPs (the larger the spacing the smaller the LP) and on the other hand the feasibility and quality of resulting trajectories. The smaller the spacing the more danger of generating jaggedness as sketched in Fig. 7, whereas the larger the spacing the more resulting trajectories obtain piecewise-affine and thus edgy shape. The danger of generating jaggedness as sketched in Fig. 7 is an intrinsic and main disadvantage of all spatial-based optimisation problem formulations. On the other hand, in the case of typical uniform spacing for a constant $D_{s}>0$ and $D_{s,j}=D_{s},~{}\forall j=0,\dots,N-1$, there is only a single hyperparameter that can be well tuned. In practice, this can be done either by trial-and-error for a given field or by a worst-case simulation analysis for given system dynamics and actuation limits. In case of in-field path planning, where there typically are no fast-moving obstacle avoidance requirements and precision path trajectories are calculated offline, worst-case simulation analysis reduces to looping over complex Dubins reference paths, calculating spatial-based LP-tracking solutions and evaluating feasibility and tracking error for a given discretisation spacing candidate $D_{s}>0$. For the numerical experiments in the next section $D_{s}=1$m is set for all instances of both Problem 1 and 2.
@@
Second, for Problem 1 a comment about the generation of a reference path is made. The presented 5-point PWA heuristic with tip-scaling serves the objective of spraying gap avoidance. However, it may not be optimal with respect to pathlength minimisation. For example, a simpler 3-point PWA heuristic might also be used as reference path. Heuristic variations on the reference path generation step taking local field contour shape information into account are the subject of future research.
Fig. 7: Sketch illustrating the danger of jaggedness for spatial-based optimisation, where characteristically along a space coordinate a lateral offset $e_{y}(s)$ (or $e_{y,j}=e_{y}(s_{j})$ in the discretised case) is assigned. The two key tuning knobs to avoid jaggedness are (i) selection of a suitable reference path such that resulting lateral deviations are close to the reference path, and (ii) selection of a suitable discretisation spacing $D_{s,j}=s_{j+1}-s_{j},~{}\forall j=0,\dots,N-1$.
4 Results
@@
Fig. 8: Experiment: 3 instances of Problem 1 and 16 instances of Problem 2 needed to be solved. The path obtained from proposed method is visualised. The sequence of nodes, $[0,16,\dots]$, gives rise to a full field coverage path plan and is calculated by the routing algorithm from Plessen (2019b) subject to the constraint that a full initial headland coverage (from node 00 to node 00 along the headland path before servicing the mainfield lanes) is sought as part of the full area coverage path. The field entrance is denoted by node 00. The inter-lane distance is 20m.
Fig. 9: Illustration of a Bezier curve-fit of third order.
@@
The solution of proposed method is presented in Fig. 8. Parameters used throughout are stated in Table 1.
Fig. 9 shows a zoom-in for an alternative solution approach to Problem 1. Here, a Bezier-curve (Pastva (1998)) of third order is fitted to the 5-point PWA path. A benefit is its simplicity, requiring only the solution of a least-squares problem to identify parameters before the evaluation of a Bernstein polynomial. A disadvantage is that eventhough the interpolating fit is smooth this method does not take vehicle actuation limits into account. The path is therefore in general not feasible to drive. In contrast, proposed LPs (6) and (7) do account for vehicle actuation limits.
Solve times and sizes of LPs for Problem 1 and 2 are in Tables 2 and 3, respectively. Sizes of the LPs are typically larger for Problem 1, with a larger average number of constraints $\bar{n}_{\text{cstrts}}$ and a larger average number of variables $\bar{n}_{\text{u}}$. Solve times are very small, which is encouraging for the general approach. All experiments were run on a laptop running Ubuntu 22.04 equipped with an Intel Core i9 CPU @5.50GHz×32 and 32 GB of memory. For the solution of the LPs Scipy’s (cf. Virtanen et al. (2020)) linprog()-solver in default settings was used.
Fig. 10 illustrates a zoom-in for sprayed coverage area avoiding any area coverage gaps.
In view of precision agriculture and autonomous robot applications an empirical finding of this work was that planning of transitions between headland path and mainfield lanes based only on Dubins-paths is not a viable option. Signification closed-loop tracking errors would be expected. As shown in Fig. 6(a) there are significant deviations between Dubins path and the path that is actually feasible considering actuation rate limits, especially when using the minimal vehicle turning radius, $R_{\text{Dubins}}=5\text{m}$, for Dubins planning .
For three different turning radii, $R_{\text{Dubins}}\in\{5\text{m},5.33\text{m},7\text{m}\}$, used in the reference path generation step (Step 1 of the hierarchical two-step algorithm), Table 4 illustrates the effect on lateral deviations between the Dubins-path used for reference path generation and the final smoothed path resulting from the LP-solution when accounting for vehicle actuation limits. The effect of using different $R_{\text{Dubins}}$ is significant. See also Fig. 6 for visualisation.
@@ 5 Discussion
A general comment about importance of input data quality is made. The quality of results for data-dependent techniques is a function of input data. For very edgy field contour input data, as is the case for all 5 fields in Fig. 8 and 12, a decision has to be made: (i) achieving full area coverage but exceeding contours when planning a nonholonomic path (see Fig. 9), or (ii) permitting area coverage gaps but achieving a path with sprayed area fully inside field contours. Both objectives cannot be achieved simultaneously. According to the formulations of Problem 1 and 2 avoiding area coverage gaps is the priority of this paper. Objective (ii) is likewise not trivial and subject of future work.
By counter-examples and discussion, below section is meant to provide further supporting arguments for proposed method.
In Oh et al. (2023) a semi-analytic path planner using three consecutive clothoids is proposed. Using continuity conditions and boundary conditions for three clothoids 8 nonlinear algebraic equations with 15 variables are derived. Using algebraic manipulations these can be reduced to two nonlinear equations for the two unknowns, which can be solved using Newton’s method and involving two tuning parameters. This method is interesting because of its analytic approach. However, it has three drawbacks that render it not suitable for our application. The first is that it is based on only three clothoids. This, however, is not sufficient to approximate all six combinations of Dubins paths (Dubins (1957)). To see this by counter-example, consider a RSR-combination (right turning, straight driving, right turning). The second disadvantage is that it is not straightforward how to adequately set the rate of curvature as a tuning parameter. For example, for Problem Type 1 steering is required that is not always fully operating at the actuating limits. To make a semi-analytic path planner work (after any extension of the approach to all 6 Dubins combinations), in such a scenario heuristics have to be devised and the method thereby evolves into a sampling scheme. The third limitation is that this analytic approach is limited to a specific vehicle model, namely one with no slip conditions at the front and the rear wheels and the center of gravity being located at the center of the rear axle.
In contrast, when extending the kinematic bicycle model (Rajamani (2011)) with slip to the model
@@
where $\beta=\text{arctan}\left(\frac{\delta}{2}\right)$ denotes the slip angle, $a$ denotes acceleration control, and the center of gravity is now located in the middle at distance $\frac{l}{2}$ to both front and rear wheels, then the same conceptual approach described in this paper can be applied. The only difference is replacement of (1) with (8). See also Miao et al. (2022) for an alternative vehicle model. Similarly, trailer dynamics can be incorporated (model equations are e.g. in Oksanen & Visala (2004), Plessen & Bemporad (2017)) into the proposed methodology.
@@
Following similarly to Oh et al. (2023) an analytic approach and as earlier work within the agricultural context, Sabelhaus et al. (2013) is briefly discussed. They present analytic formulas and visualisations for 9 different symmetric lane-to-lane transitions (Omega-turn, U-turn, Gap-turn, Slope-turn, Fishtail-turn, Fishtail-turn with slope, Minimal longitudinal width, Pinhole-turn, Reversal pinhole turn) for continuous-curvature path planning. They also present three simulation experiments with longitudinal, lateral and angle offset of start and end point. Characteristically, all paths are constructed based on path segments with maximum curvature rate (clothoids with so called maximal sharpness) and circle segments with maximum curvature. Five comments are made. First, the methods are not suitable for Problem 1 since for that problem paths are needed that are not always and necessarily operating at the maximum curvature rate limit. Second and interestingly, in general none of the 9 cases can handle Problem 2 (e.g. for transitions as shown in Fig. 6). The Minimal longitudinal width-case resembles a transition as shown in Fig. 6, but it has a specific condition $d=R_{\text{big}}$ (Sect. 2.3.7 in Sabelhaus et al. (2013)) that limits its usage. (Another reason herefore is that Sabelhaus et al. (2013) focus only on lane-to-lane transitions and not dicuss headland-to-lane transitions.) Third, Sabelhaus et al. (2013) acknowledge that ’manoeuvres with start or end curvature $\kappa\neq 0$ are not feasible with continuous curvature turns’. In contrast, for proposed LP-based method this case can be handled, by adding an initial input constraint $\delta_{0}=\delta_{\text{start}}$ or $\delta_{N-1}=\delta_{\text{end}}$ to (6) or (7) for Problem 1 or 2, respectively, and by observing that curvature can be related to steering via $\kappa=\frac{\text{tan}(\delta)}{l}$ (for the assumption of no-slip conditions of the wheels and instantaneous curvature of the rear axle). This is made possible since the LP-based approach permits to incorporate state and input constraints in a structured manner. Fourth, as discussed for Oh et al. (2023), likewise the approach of Sabelhaus et al. (2013) is limited to a specific vehicle model (no slip condition). Fifth, the proposed spatial LP-method can also handle lane-to-lane transitions. For illustration see Fig. 11 with multiple non-symmetric start and end points and an interlane-distance of 5m. The simplicity of proposed method is underlined. For lane-to-lane transitions a Dubins-path is computed between the intersection points of headland path and mainfield lanes. The Dubins path is then used as reference path in the second hierarchical step solving a spatial-based LP to obtain a refined path. This paragraph underlines the flexibility of proposed method, in particular for shaping solutions to desired behavior.

### v4

@@ 3.3 Problem 1: Headland path smoothing
In a first step, the segment is scaled by heuristics and replaced by a 5-point piecewise-affine (PWA) path as shown in Fig. 4(a) under certain conditions. First, the key characteristic of this scaling is that the tip of the resulting 5-point PWA path shall ensure zero spraying gap. Thus, the distance between the tip and its projection point along the field contour must be less or equal than half the machinery operating width $w$. Secondly, beyond this key characteristic, the shaping of the reference path is a heuristic and free choice. Here, a 5-point PWA path segment is used because it rounds out the original edgy headland path segment (more than, for example, an alternative 3-point PWA path segment would) and thereby helps the downstream linear program formulation by providing an improved reference around which the system dynamics are linerised and discretised. Thirdly, the heuristic nature of this first reference path generation step is acknowledged and a potential avenue of future research. Lastly, and as a detail, the popular Chaikin’s algorithm (see Chaikin (1974)) is not used for corner cutting and further smoothing of the reference path, since this would replace every edge point with two edge points at each iteration step. Furthermore, it would cut off a significant portion of the tip of the 5-point PWA line after the first iteration.
In a second step, using the 5-point PWA path from Step 1 as a reference path, the following basic LP is proposed:
@@
Steering actuation constraints are in Eqn. (6c) and (6d) as discussed in Sect. 3.1. In practice, constraints in Eqn. (6b), which enforces the resulting path to stay on one side of the reference path, is relaxed to, $e_{y,j}-e_{y,j}^{\text{ref}}\leq\sigma$, by introducing a non-negative slack variable $\sigma\geq 0$ and adding it to the objective function 6a with a high weight such that $\sum\nolimits_{j=1}^{N}|e_{y,j}-e_{y,j}^{\text{ref}}|$ is replaced with $\sum\nolimits_{j=1}^{N}|e_{y,j}-e_{y,j}^{\text{ref}}|+10^{16}\sigma$. Furthermore, for the absolute value arguments, $N$ non-negative surrogate variables are introduced and the set of inequality constraints is extended accordingly. In the end, there are $n_{u}=2N+1$ scalar real-valued optimisation variables. Typical sizes of the LP (i.e., the number of variables and inequality constraints) are stated in the numerical example Section 5.
For generality, reference $e_{y,j}^{\text{ref}}$ is included in Eqn. (6a) and (6b). In practice, it is typically set to zero, $e_{y,j}^{\text{ref}}=0,\forall j=1,\dots,N$.
One benefit of proposed formulation in LP (6) is that it is hyperparameter-free. There are no weighting parameters.
An even smaller alternative LP was considered, with a minmax-objective function, $\min\nolimits_{\{\delta_{j}\}_{j=0}^{N-1}}\ \max\nolimits_{j\in\{1,\dots,N\}}|e_{y,j}-e_{y,j}^{\text{ref}}|$, replacing Eqn. (6a). However, this was found to be unsuited since only the maximum deviation from the reference path would be penalised, which resulted in jagged trajectories not accurately tracking the complete reference path.
After the solution of LP (6), a smoothed headland path segment as shown in Fig. 4(b) results.
Finally, LP-iterations or refinement steps can be conducted. Thus, for the first solution of LP (6) a 5-point-PWA path is used as a reference. Then, the resulting path is used as a reference for the second solution of LP (6). In theory, many such LP-iterations can be carried out. In practice, one refinement step was found to be useful and sufficient.
@@ 3.4 Problem 2: Headland-to-mainfield lanes transitions
Before stating the LP for Problem 2, the term headland-to-mainfield lanes transitions is clarified. It is used comprehensively for brevity but shall always refer to both, (i) transitions from a headland to a mainfield lane, but also (ii) transitions from a mainfield lane to the headland path. Both types of transitions naturally occur for a field coverage path. As will be shown below, a subtle difference in constraints formulation is applied to differentiate the two types.
@@
In practice, the objective function Eqn. (7a) is slightly refined to better avoid undesired overshoots out of the mainfield area and into the headland area. Therefore, Eqn. (7a) is replaced with $\sum\nolimits_{j=1}^{N}c_{j}|e_{y,j}-e_{y,j}^{\text{ref}}|$ and weights
@@
for transitions from mainfield lane to the headland path, where index $I$ is computed from the reference path as its last and first index within a 1 m radius to the headland path for the two transition types. The value 100 is selected exemplarily and can be any large constant, such that the objective function of the LP can still be considered hyperparameter-free. The benefit of this weighted approach is that no additional inequality constraints need to be introduced. As a result, LP (7) for Problem 2 is smaller than LP (6) for Problem 1.
As for above LP (6), for the absolute value arguments in Eqn. (7a), $N$ non-negative surrogate variables are introduced and the set of inequality constraifnts is extended accordingly. In the end, there are $n_{u}=2N$ scalar real-valued optimisation variables. Typical sizes of the LP (i.e., the number of variables and inequality constraints) are stated in the numerical example in Section 5. As in the previous subsection, reference $e_{y,j}^{\text{ref}}$ in LP (7) is typically set to zero, $e_{y,j}^{\text{ref}}=0,\forall j=1,\dots,N$.
Figure 6 visualises the result of a headland-to-mainfield lane transition. As part of Step 1 of the hierarchical algorithm for Problem 2, a Dubins path is fitted and replaces the edgy original transition. The Dubins path, plus an additional small path segment before and after the Dubins path to permit additional maneuvering space, serves as reference to Step 2 of the hierarchical algorithm based on which LP (7) is formulated and solved.
Figures 5 and 6 are of special interest for two reasons. First, they illustrate how Dubins paths are not suited for the generation of precision paths to be followed by autonomous machines online, especially when calculating Dubins paths with minimal turning radius of the machine (see Fig. 6(a)). Due to actuation constraints, deviations would be expected in any online tracking application. Second, it illustrates the effect on tracking performance when using different turning radii $R_{\text{Dubins}}$ in the Dubins reference path generation step. When using the minimal turning radius, results in the deviation shown in Fig. 6(a) result. By using an artificially larger $R_{\text{Dubins}}$, the LP-solution trajectory more closely follows the reference path, see Fig. 6(b) and 6(c). The drawback of a larger $R_{\text{Dubins}}$ is that the resulting path is longer and a larger area of compaction will occur from corresponding tyre traces and the corresponding steering maneuvers, which reduces the productive crop area in the field. A trade-off with a preference towards small $R_{\text{Dubins}}$ for the reference path generation step is recommended.
Finally, in contrast to Problem 1 and its discussion in Section 3.3 for the case of Problem 2, no LP-iterations are conducted. It was empirically found that they did not improve results and are omitted.
@@
Fig. 6: Problem 2: The effect on tracking performance when using three different $R_{\text{Dubins}}$ in the reference path generation step is visualised. Paths denoted with ’Reference’ and ’LP spatial’ denote the paths before and after the LP-solutions, respectively.
4 Methods
@@
This section elaborates on important parameter choices and other practical details. First, the selection of discretisation spacing, $D_{s,j}=s_{j+1}-s_{j},~{}\forall j=0,\dots,N-1$, is discussed, which typically is selected as uniformly spaced but can also be non-uniformly spaced. Its importance for spatial-based numerical optimisation is underlined. Its selection influences both the size of the LPs (the larger the spacing the smaller the LP) and the feasibility and quality of resulting trajectories. The smaller the spacing, the more danger there is of generating jaggedness as sketched in Fig. 7, whereas the larger the spacing the more the resulting trajectories obtain piecewise-affine and thus an edgy shape. The danger of generating jaggedness, as sketched in Fig. 7, is a disadvantage of the spatial-based optimisation problem formulation. On the other hand, in the case of typical uniform spacing for a constant $D_{s}>0$ and $D_{s,j}=D_{s},~{}\forall j=0,\dots,N-1$, there is only a single hyperparameter that can be well tuned. In practice, this can be done either by trial-and-error for a given field or by a worst-case simulation analysis for given system dynamics and actuation limits. In case of in-field path planning, where there typically are no fast-moving obstacle avoidance requirements and precision path trajectories are calculated offline, worst-case simulation analysis reduces to looping over complex Dubins reference paths, calculating spatial-based LP-tracking solutions and evaluating feasibility and tracking error for a given discretisation spacing candidate $D_{s}>0$. For the numerical experiments in the next section $D_{s}=1$ m is set for all instances of both Problem 1 and 2.
@@
Second, for Problem 1 a comment about the generation of a reference path is made. The presented 5-point PWA heuristic with tip-scaling serves the objective of spraying gap avoidance. However, it may not be optimal with respect to pathlength minimisation. For example, a simpler 3-point PWA heuristic might also be used for a reference path. Heuristic variations on the reference path generation step taking local field contour shape information into account will be the subject of future research.
All experiments were run on a laptop running Ubuntu 22.04 equipped with an Intel Core i9 CPU @5.50GHz×32 and 32 GB of memory. For the solution of the LPs Scipy’s (cf. Virtanen et al. (2020)) linprog()-solver in default settings was used.
Fig. 7: Sketch illustrating the danger of jaggedness for spatial-based optimisation. A reference path (black) is desired such that lateral deviations resulting from the LP-solution (blue) remain small. This is the case in the right drawing, but not in the left drawing.
@@
5 Results
Fig. 8: Experiment: 3 instances of Problem 1 and 16 instances of Problem 2 needed to be solved. The path obtained from proposed method is visualised. The sequence of nodes, $[0,16,\dots]$, gives rise to a full field coverage path plan and is calculated by the routing algorithm from Plessen (2019b) subject to the constraint that a full initial headland coverage (from node 00 to node 00 along the headland path before servicing the mainfield lanes) is sought as part of the full area coverage path. The field entrance is denoted by node 00. The inter-lane distance is 20 m. Tracing the number sequence in the heading gives rise to the area coverage path.
Fig. 9: Illustration of a zoom-in for an alternative solution approach to Problem 1: a Bezier curve of third order is fitted to the 5-point PWA path. This approach does not explicity account for vehicle actuation constraints. For comparison, see the bottom turn in Fig. 8 and 10, which results from the solution of LP (6) and takes vehicle actuation constraints into account.
@@ 5 Results
The solution of proposed method is presented in Fig. 8. Parameters used throughout are stated in Table 1. Fig. 9 shows a zoom-in for an alternative solution approach to Problem 1. Here, a Bezier-curve (Pastva (1998)) of third order is fitted to the 5-point PWA path. This implies solving a least-squares problem to identify parameters before the evaluation of a Bernstein polynomial to generate the Bezier curve-fit. A benefit is the simplicity of this approach. A disadvantage is that even though the interpolating fit is smooth, this method does not take vehicle actuation limits into account. Therefore, the path is in general not feasible to drive. In contrast, proposed LPs (6) and (7) do account for vehicle actuation limits.
The solution times and sizes of LPs for Problem 1 and 2 are in Tables 2 and 3, respectively. Sizes of the LPs are typically larger for Problem 1, with a larger average number of constraints ($\bar{n}_{\text{cstrts}}$) and a larger average number of variables ($\bar{n}_{\text{u}}$). The solution times are very small, which is encouraging for the general approach.
[…]
@@ 6 Discussion
In view of precision agriculture and autonomous robot applications an empirical finding of this work was that planning of transitions between headland path and mainfield lanes based only on Dubins-paths was not a viable option. Signification closed-loop tracking errors would be expected. As shown in Fig. 6(a) there are significant deviations between Dubins path and the path that is actually feasible considering actuation rate limits, especially when using the minimal vehicle turning radius, $R_{\text{Dubins}}=5\text{~{}m}$, for Dubins planning.
A general comment about importance of input data quality is made. The quality of results for data-dependent techniques is a function of input data. For very edgy field contour input data, as is the case for all 5 fields in Fig. 8 and 11, a decision has to be made: (i) achieving full area coverage but exceeding contours when planning a nonholonomic path (see Fig. 9), or (ii) permitting area coverage gaps but achieving a path with sprayed area fully inside field contours. Both objectives cannot be achieved simultaneously. According to the formulations of Problem 1 and 2 avoiding area coverage gaps is the priority of this paper. Objective (ii) is likewise not trivial and a subject for future work.
By counter-examples and discussion, the following provides further supporting arguments for the proposed method.
In Oh et al. (2024) a semi-analytic path planner using three consecutive clothoids was proposed. Using continuity conditions and boundary conditions for three clothoids, 8 nonlinear algebraic equations with 15 variables were derived. Using algebraic manipulations these can be reduced to two nonlinear equations for the two unknowns, which can be solved using Newton’s method involving two tuning parameters. This method is interesting because of its analytic approach. However, it has three drawbacks that render it unsuitable for the applications in this paper. The first is that it is based on only three clothoids, which is insufficient to approximate all six combinations of Dubins paths (Dubins (1957)). To see this by counter-example, consider a RSR-combination (right turning, straight driving, right turning), which requires 4 clothoids for path smoothing. The second disadvantage is that it is not straightforward to correctly set the rate of curvature as a tuning parameter. For example, for Problem Type 1, steering is required that is not always fully operating at the actuating limits. To make a semi-analytic path planner work (after any extension of the approach to all 6 Dubins combinations), scenario heuristics have to be devised and the method thereby evolves into a sampling scheme. The third limitation is that this analytic approach is limited to a specific vehicle model, namely one with no slip conditions at the front and the rear wheels and the centre of gravity being located at the centre of the rear axle.
In contrast, when extending the kinematic bicycle model (Rajamani (2011)) with slip to
@@
where $\beta=\text{arctan}\left(\frac{\delta}{2}\right)$ denotes the slip angle, $a$ denotes acceleration control, and the centre of gravity is now located in the middle at distance $\frac{l}{2}$ to both front and rear wheels, then the same conceptual approach described in this paper can be applied. The only difference is the replacement of Eqn. (1) with (8). See also Miao et al. (2022) for an alternative vehicle model. Similarly, trailer dynamics can be incorporated (equations in Oksanen & Visala (2004), Plessen & Bemporad (2017)) into the proposed methodology.
@@
Similarly to Oh et al. (2024), Sabelhaus et al. (2013) applied an analytic approach in earlier work in an agricultural context. They presented analytic formulas and visualisations for 9 different symmetric lane-to-lane transitions (Omega-turn, U-turn, Gap-turn, Slope-turn, Fishtail-turn, Fishtail-turn with slope, Minimal longitudinal width, Pinhole-turn, Reversal pinhole turn) for continuous-curvature path planning. They also presented three simulation experiments with longitudinal, lateral and angle offset of start and end point. Characteristically, all paths were constructed based on path segments with maximum curvature rate (clothoids with so called maximal sharpness) and circle segments with maximum curvature. Five comments are made. First, the methods are not suitable for Problem 1 as problem paths are needed that are not always and necessarily operating at the maximum curvature rate limit. Second and interestingly, in general, none of the 9 cases can handle Problem 2 (e.g. for transitions as shown in Fig. 6). The Minimal longitudinal width-case resembles a transition as shown in Fig. 6, but it has a specific condition $d=R_{\text{big}}$ (Sect. 2.3.7 in Sabelhaus et al. (2013)) that limits its usage. Another reason is that Sabelhaus et al. (2013) focussed only on lane-to-lane transitions and did not dicuss headland-to-lane transitions. Third, Sabelhaus et al. (2013) acknowledge that ’manoeuvres with start or end curvature $\kappa\neq 0$ are not feasible with continuous curvature turns’. In contrast, the proposed LP-based method this case can be handled, by adding an initial input constraint $\delta_{0}=\delta_{\text{start}}$ or $\delta_{N-1}=\delta_{\text{end}}$ to LP (6) or (7) for Problem 1 or 2, respectively, and by observing that curvature can be related to steering via $\kappa=\frac{\text{tan}(\delta)}{l}$ (for the assumption of no-slip conditions of the wheels and instantaneous curvature of the rear axle). This is made possible as the LP-based approach incorporated state and input constraints in a structured manner. Fourth, as discussed for Oh et al. (2024), the approach of Sabelhaus et al. (2013) is limited to a specific vehicle model (no slip condition).
The proposed spatial LP-method can also handle lane-to-lane transitions. For illustration see Fig. 12 with multiple non-symmetric start and end points and an interlane-distance of 5 m. The simplicity of the proposed method is that for the lane-to-lane transitions a Dubins-path is computed between the intersection points of headland path and mainfield lanes. The Dubins path is then used as the reference path in the second hierarchical step solving a spatial-based LP to obtain a refined path. This underlines the flexibility of the proposed method, in particular for shaping solutions to desired behavior.